Structuring a Keep-or-Drop Product Line Problem with Complementary Effects

Shown below is a segmented income statement for Hickory Company's three wooden flooring product lines:

Strip Plank Parquet Total
Sales revenue $400,000 $200,000 $300,000 $900,000
Less: Variable expenses 225,000 120,000 250,000 595,000
Contribution margin $175,000 $ 80,000 $ 50,000 $305,000
Less direct fixed expenses:
   Machine rent (5,000) (20,000) (30,000) (55,000)
   Supervision (15,000) (10,000) (5,000) (30,000)
   Depreciation (35,000) (10,000) (25,000) (70,000)
Segment margin $120,000 $ 40,000 $ (10,000) $150,000

Hickory's management is deciding whether to keep or drop the parquet product line. Hickory's parquet flooring product line has a contribution margin of $50,000 (sales of $300,000 less total variable costs of $250,000). All variable costs are relevant.

Relevant fixed costs associated with this line include 80% of parquet's machine rent and all of parquet's supervision salaries. In addition, assume that dropping the parquet product line would reduce sales of the strip line by 22% and sales of the plank line by 20%. All other information remains the same.

Required:

1. If the parquet product line is dropped, what is the contribution margin for the strip line?
$

For the plank line?
$

2. Which alternative (keep or drop the parquet product line) is now more cost effective and by how much?
  by $
Answer #1

Solution

Hickory Company

Analysis of alternative – Drop parquet product line which would reduce the sales of strip line by 22% and sales of plank line by 20%:

Revised sales of strip line – 400,000 – 22% x 400,000 = $312,000

Revised sales of plank line – 200,000 – 20% x 200,000 = $160,000

Contribution format income statement:

Hickory Company

Contribution margin income statement

Strip

Plank

Parquet

Total

Sales

$312,000

$160,000

0

$472,000

Less: Variable expenses

$175,500

$96,000

0

$271,500

Contribution margin income statement

$136,500

$64,000

0

$200,500

Less: Direct fixed expenses

machine rent

$5,000

$20,000

6,000

$31,000

Supervision

$15,000

$10,000

$25,000

Depreciation

$35,000

$10,000

$25,000

$70,000

Total direct expenses

$55,000

$40,000

$31,000

$126,000

Segment margin

$81,500

$24,000

($31,000)

$74,500

The contribution margin for the strip line is $136,500.

The contribution margin for the plank line is $64,000.

Note:

The unavoidable machine rent for parquet line = 30,000 x 20% = $6,000

The depreciation expense of parquet line is unavoidable.

The variable expenses are computed based on the original sales.

  • Strip line = 312,000 x 225,000/400,000 = $175,500
  • Plank line = 160,000 x 120,000/200,000 = $96,000

The resulting decrease in sales percentages for strip line and plank line due to dropping of parquet product line caused a reduction in contribution margin of strip line and parquet line.

Also, the company as a whole would lose the contribution margin of parquet product line, $50,000.

The combined segment margin is $74,500, which is less by $75,500 from the original segment margin of $150,000. Hence, dropping parquet line will drop segment margin by $75,500.

Hence, the cost effective decision is to keep parquet product line.